# Design of a LYSO Crystal Electromagnetic Calorimeter for DarkSHINE Experiment

### v1

@@ Abstract
This paper presents the design and optimization of a LYSO crystal-based electromagnetic calorimeter (ECAL) for the DarkSHINE experiment, which aims to search for dark photon as potential dark force mediator. The ECAL design has been meticulously evaluated through comprehensive simulations, focusing on optimizing dimensions, material choices, and placement within the detector array to enhance sensitivity in search for dark photon signatures while balancing cost and performance. The concluded ECAL design, comprising 2.5$\times$2.5$\times$4 cm3 LYSO crystals arranged in a 52.5$\times$52.5$\times$44 cm3 structure, ensures high energy resolution and effective energy containment. The study also explored the energy distribution across different ECAL regions and established a dynamic range for energy measurements, with a 4 GeV limit per crystal deemed sufficient. Additionally, the radiation tolerance of ECAL components was assessed, confirming the sustainability of LYSO crystals and radiation-resistant silicon sensors.
@@ I Introduction
Dark matter (DM) Hooper:2007kb ; Clowe_2006 remains one of the most compelling mysteries in cosmology and particle physics, evident from its gravitational effects on visible matter and the cosmic microwave background. The prevailing hypothesis is that DM, which does not interact with ordinary matters, constitutes about 85% of the universe’s total mass. Understanding DM is crucial for unraveling the universe’s structure and origins.
@@
Theoretical models suggest that DM was produced through thermal processes in the early universe, with the "freeze-out" mechanismDu:2021jcj explaining its current observed density, positing a probable mass range from a few MeV to several TeV. Despite extensive searches, the specific properties and particle nature of DM remain elusive. Experiments such as XENONnTXENON:2023cxc , PandaXPandaX-4T:2021bab , CDEXWang_2021 , LUX-ZEPLINPhysRevLett.131.041002 , AMSGiovacchini:2020vxz , DAMPEKyratzis:2022gvg and LHCGiagu:2019fmp have narrowed the parameter space for weakly interacting massive particles (WIMPs)Griest:1989wd ; Ho:2012ug ; Steigman:2013yua ; Boehm:2013jpa ; Nollett:2013pwa ; Nollett:2014lwa ; Serpico:2004nm in the GeV to TeV range, yet these particles remain undetectedBillard:2021uyg .
@@
Given the high-mass DM candidate particles are yet to be detected, there is increasing interest in the sub-GeV mass range, which presents significant detection challenges due to the minimal interaction cross-sections of such light particles with ordinary matter. This has necessitated the development of innovative detection techniques distinct from traditional DM detection methods. Accelerator-based experiments are particularly promising, leveraging high-energy particle collisions to produce and detect dark photons, a hypothetical mediator between visible and dark matterHoldom:1985ag ; Foot:1991kb ; Fuyuto:2019vfe ; Choi:2020pyy ; Cheng_2022 . Facilities such as CERN’s NA64Banerjee:2019pds , LHC2024138378 , BELLE-IIPhysRevLett.130.071804 , BES-IIIPhysRevD.100.115016 ; Prasad:2019ris , and the proposed LDMXLDMX experiment aim to explore these possibilities.
@@
Figure 1: Panel (a) Production of dark photons via bremsstrahlung. (b) Decay of dark photons into "invisible" modes, where the dark photon decays into dark matter particles that do not interact with ordinary matter. Fabbrichesi:2020wbt
@@
The DarkSHINE experimentDarkSHINE ; SLi , designed to operate under the minimal dark photon model, proposes a novel approach to detecting these elusive particles. Figure 1 illustrates the conceptual framework of dark photon production through dark bremsstrahlung and their subsequent decay into invisible dark matter, forming the experimental basis for DarkSHINE. This experiment leverages the high repetition rate electron beams provided by the Shanghai High Repetition-Rate XFEL and Extreme Light Facility (SHINE)Wan:2022het ; Zhao:2017ood , optimized for detecting subtle signals that hint the presence of dark photons.
@@
This paper delineates the design principles of the Electromagnetic Calorimeter (ECAL), focusing on the selection of materials and structural layout. Through comprehensive simulation studies, we have optimized the ECAL configuration, achieving a balance between performance and cost-efficiency. We also explored the dynamic range of energy that the ECAL’s crystals can absorb. Insights from these analyses have guided the development of potential triggering strategies for future detectors. Additionally, we assessed the radiation tolerance of the ECAL, highlighting the necessity for crystals and silicon sensors that maintain high performance in high-radiation environments.
@@ II DarkSHINE experiment and ECAL design
DarkSHINE is a fixed-target experiment that focuses on the bremsstrahlung production of dark photons and measuring their invisible decay. It utilizes a high repetition rate single-electron beam provided by SHINE, which is currently under construction. The designed beam energy from SHINE is 8 GeV with a designed repetition rate of 1 MHz. The beam is expected to achieve an energy of 8GeV and a repetition rate of 10MHz. This corresponds to $3\times 10^{14}$ electron-on-target (EOT) events during one year of the DarkSHINE experiment’s commissioning.
Figure 2: Schematic of the DarkSHINE detector system, illustrating the primary components used for dark photon detection. The setup includes a tungsten target, a tagging tracker, a recoil tracker, an electromagnetic calorimeter (ECAL), and a hadronic calorimeter (HCAL)DarkSHINE .
@@
The primary challenge for DarkSHINE is to use information from various detectors to minimize background contributions while preserving the dark photon signal. Figure 2 shows the detector system of DarkSHINE. It primarily consists of a tungsten target (0.1 X0) and three sub-detectors: a tracker, an electromagnetic calorimeter (ECAL), and a hadronic calorimeter (HCAL). Additionally, a single electron beam with a spatial distribution radius of 3 cm and a maximum non-uniform magnetic field of 1.5 Tesla in the tracker region were used in the simulations. The tracker measures the momentum and trajectory of electrons, consisting of a total of thirteen layers of silicon strips. Seven layers are placed in front of the target to tag the incident electron (tagging tracker), and six layers are placed behind the target (recoil tracker). A crystal ECAL positioned behind the recoil tracker measures the deposited energy of recoil electrons and photons. Then, a scintillator-steel based sampling HCAL is placed behind and surrounding the ECAL to capture and veto the backgrounds, in particular the neutral hadrons and muons.
@@
To measure the invisible decay of dark photons, it is essential to precisely measure the energy of recoil electrons after each collision to determine if there is a significant energy loss. The core detector of DarkSHINE, the ECAL designed for DarkSHINE, is a homogeneous LYSO crystal calorimeter with very high electromagnetic energy resolution. LYSO1239590 ; CMS:2667167 ; Kalinnikov:2023coj is an excellent choice for our crystal material, not only because of its high light yield, which is critical for electromagnetic resolution, but also due to its rapid scintillation decay time (40 ns), which is vital for handling the extremely high event rate we face. Furthermore, the ECAL, especially its central region, is exposed to a considerable radiation dose, necessitating materials with good radiation resistance. LYSO has shown exceptional durability in such conditions, establishing it as an optimal selection for DarkSHINE ECAL applications.
The radiation length of LYSO is about 1.12 cm. The transverse dimension of the ECAL is 52.5$\times$52.5 cm3, with a dimension along the beam direction of 44 cm, approximately equivalent to 39 radiation lengths. The substantial depth ensures excellent energy containment in the ECAL, guaranteeing that it can absorb nearly all electromagnetic showers, preventing them from leaking into the HCAL. This is crucial because such leakage would result in vetoing dark photon signals by the HCAL. The ECAL is fully segmented to gather comprehensive information, enhancing its capability to discriminate dark photons. After optimization, the ECAL is composed of 21$\times$21$\times$11 LYSO crystals, with each crystal measuring 2.5$\times$2.5$\times$4 cm3 and read out by SiPMsDINU2016255 ; SIMON201985 . In each layer, the crystals are positioned in a uniform 21$\times$21 square pattern. To enhance detection efficacy and prevent particles from traversing the gaps, the placement of crystals in successive layers is shifted by half the transverse dimension of a crystal. This staggered structure aids in detecting recoil electron with energy loss coming along with dark photon signal.
III Simulation, digitization and optimization
III.1 Simulation and digitization
III.1.1 Software setup
@@
Figure 3: Configuration of the Electromagnetic Calorimeter (ECAL) in DarkSHINE. The ECAL consists of an array of LYSO crystals, each coupled with SiPMs for readout. The schematic shows the segmentation and staggered layout of the crystals within the ECAL.
@@
The Monte Carlo simulation of the ECAL is performed using the DarkSHINE Software (DSS) framework based on Geant4GEANT4:2002zbu . This simulation includes the entire detector setup, encompassing the crystal cubes, wrappers, SiPM sensors, and support structures, alongside other components like the Tracker and HCAL. This comprehensive setup provides the necessary reference for ECAL reconstruction and the initial analysis cuts to define the appropriate phase space for ECAL analysis. The detector configuration used in the simulation is illustrated in Figure 3. Within DSS, we can simulate inclusive events that include all background processes or individual simulations of dark photon processes and specific rare background events.
@@
III.1.2 Energy Digitization
Accurate simulation of the response of crystal-based detectors requires a precise description of digitization effects. Digitization applies a series of realistic effects to the simulation results, mimicking the behavior of an actual detector and bringing the simulation more comparable to real experimental conditions. The energy directly obtained from Geant4, known as the truth energy, represents the ideal energy deposition of a particle in a perfect detection scenario, devoid of any detector effects or measurement errors. Thus, digitization provides a more realistic depiction of detector performance, offering stronger evidence for optimization.
The main goal of energy digitization is to parameterize the behavior of each element related to energy measurement. This involves applying smearing based on experimental data. For the DarkSHINE ECAL, the energy digitization process can be primarily divided into three parts: scintillation digitization, SiPM digitization, and ADC digitization.
Table 1: Parameters in scintillation digitization.
Scintillator Light yield (intrinsic) Light yield fluctuation
LYSO 30000 ph/MeV 10%
Light yield calibration accuracy SiPM PDE (average) Light yield (measured)
1% 6.3% 150 p.e./MeV
The first step is scintillation digitization. We set an intrinsic light yield of 30,000 photons per MeV for the LYSO crystal scintillator, assuming a 10% fluctuation across all scintillators. The scintillation light is proportional to the energy deposited in the crystal and follows a Poisson distribution. The scintillation light experiences attenuation during propagation within the crystal and detection by the SiPM. The attenuation depends on the photon transportation length and the detection efficiency of the SiPM. The measured light yield helps to determine the amount of light loss during propagation. Although the measured light yield for all crystal-SiPM units shows fluctuations, these can be calibrated to some extent, assumed here to have a 1% accuracy. The parameters used in this step are shown in Table 1.
Figure 4: Simulated SiPM response for HAMAMATSU S14160-3010PS, illustrating the relationship between the number of incident photons ($N_{In}$) and the number of fired pixels ($N_{fired}$). This simulation takes into account photon detection efficiency (PDE), pixel density, pixel recovery time, and crosstalk effects. The simulated points are fitted with a formula (blue dashed line), providing a description of the SiPM response and a method for correcting saturation effects.
Table 2: Parameters in SiPM digitization.
Pixel number Crosstalk Gain fluctuation Gain calibration accuracy
89984 0.5% 10% 1%
For SiPM digitization, a toy Monte Carlo model simulates the SiPM’s response to the scintillation light from the LYSO crystal. This simulation is based on the HAMAMATSU S14160-3010PSS14160-3010PS , featuring a 3×3 mm2 sensitive area, 10 $\mu$m pixels, and nearly 90,000 pixels in total. This model considers the photon detection efficiency, pixel density, pixel recovery, and crosstalk effect of the SiPM. Through this model, we can determine the SiPM’s response to varying numbers of incident photons. As shown in Figure 4, $N_{fired}$ is the number of photons detected by the SiPM, $\epsilon$ is the photon detection efficiency, and $N_{In}$ is the number of photons incident on the SiPM. As the number of incident photons increases, the number of detected photons deviates from linearity and tends to saturate. A formula from kotera2016describing is used to fit the SiPM’s response. This formula helps in simulating the SiPM’s response to variable numbers of incident photons, providing a means to correct for saturation effects. Parameters in SiPM digitization are shown in Table 2.PIEMONTE20192 ; NAGY201444
Table 3: Parameters in ADC digitization.
Modes Gain ADC/p.e. Noise
High gain $\times$1 10 3 ADC
Middle gain $\times$45 0.2 3 ADC
Low gain $\times$2000 0.005 3 ADC
Equivalent noise energy Switching point Energy range
2 keV 4000 ADC 10-2.7 MeV
100 keV 4000 ADC 2.7-133.3 MeV
4 MeV - 133.3-5461 MeV
The behavior of the readout electronics is simulated in the ADC digitization process. Here, we assume that the charge produced by the SiPM is recorded by a 12-bit multi-channel analyzer with a total of 4096 ADC channels. There are three gain modes to expand the dynamic range: high gain, medium gain, and low gain. In high gain and medium gain modes, when the ADC exceeds 4000, it will automatically switch to the next gain level. The smaller the gain, the larger the equivalent electronic noise. Table 3 shows the parameters used in ADC digitization.
In Tables 1,  2, and  3, some parameters are obtained based on a 2.5$\times$2.5$\times$4 cm3 crystal, an optimized size discussed in section III.3. Each part’s contribution to the single-channel energy resolution is illustrated in Figure 5. We individually simulated the digitization process for each component to obtain its resolution and also simulated the process with all three components present to determine the overall resolution. At lower energies (less than 100 MeV), the resolution of the SiPM digitization is the main contributor. At higher energies, contributions from the scintillator digitization and ADC digitization become more significant.
Figure 5: Energy resolution of single channel for each digitization part and total effect. At energies below 100 MeV, the primary contributor to the resolution is the SiPM digitization. As the energy increases, the scintillator digitization and ADC digitization become the dominant factors.
It should be noted that the current digitization is only used for the optimization of the crystal unit size in section III.3. For other analyses involving the total ECAL energy, for simplicity, we only use the truth energy.
@@ III.2 Signal and background
Signal and background processes exhibit distinct characteristics within different detectors. Typically, incident electrons pass through the target with negligible energy loss. In the tagging tracker and recoil tracker, each electron leaves a single track, and the measured momentum remains consistent across both trackers. Nearly all recoil electron energy is deposited in the electromagnetic calorimeter (ECAL), with minimal signals detected by the hadronic calorimeter (HCAL). In the ECAL, normal electron events usually show a significant energy deposition, typically around 8 GeV.
Figure 6: Deposit energy in ECAL of different processes: (a) Dark photon and inclusive background processes. Inclusive background processes (grey) result in significantly higher energy deposition within the ECAL, reaching up to 8 GeV. In contrast, the energy deposition from dark photon signal processes is significantly lower, mainly concentrating around 0 GeV. (b) Dark photon and rare background processes. Only the bremsstrahlung process exhibits significantly distinct energy deposition in the ECAL compared to the signal, as its final state particles are electrons and photons. Other processes involving muons and nuclear reactions result in more energy deposition at the low-energy end and cannot be completely excluded by the ECAL.
@@
When a dark photon is produced during the electron-on-target process, most of the incident energy transfers to it, and the recoiled electron deposits its remaining energy in the ECAL. The dark photon decays into dark matter, leaving no signal in the tracker or calorimeter, causing significant energy loss in the ECAL, typically exceeding 4 GeV (Figure 6). As the dark photon mass increases, it carries more energy from the electrons, leading to larger energy losses of the recoiled electron measured in the ECAL.
Figure 7: Major background processes and their branching ratios in the DarkSHINE experiment. DarkSHINE
Figure 7 shows the major background processes and their relative rates. Background processes in the DarkSHINE experiment are categorized into two types: soft background, where the recoil electron loses less than 4 GeV, and hard background, where the energy loss exceeds 4 GeV.
Soft background processes, including normal electron events, dominate the total event production, constituting over 90%. These processes result in substantial energy deposition within the ECAL, making them easily distinguishable from dark photon signals.
In contrast, hard background processes remove most of the recoil electron’s energy. If electrons undergo electromagnetic interactions, such as hard bremsstrahlung or conversion of bremsstrahlung photons into electron-positron pairs at the target, the energy deposited in the ECAL remains a critical rejection criterion. However, processes involving photon conversion into muon pairs are not easily distinguished by the ECAL and require exclusion by the HCAL and tracker. Background processes involving nuclear reactions, such as electron-nucleus interactions or bremsstrahlung photons interacting with nuclei, produce secondary particles with hadronic components, primarily neutronswang2024design . These particles penetrate the ECAL and are to be vetoed by the HCAL or the tracker.
@@
Neutrino-producing background processes, such as Moller scattering ($e^{-}e^{-}\rightarrow e^{-}e^{-}$) followed by a charged-current quasi-elastic (CCQE) reaction ($e^{-}p\rightarrow\nu_{e}n$), neutrino pair production ($e^{-}N\rightarrow e^{-}N\nu\nu$), bremsstrahlung with CCQE, and charge-current exchange with exclusive ($e^{-}p\rightarrow\nu n\pi^{0}$), as mentioned in Izaguirre:2014bca , have been found negligible in this experiment due to their low event ratios. Therefore, these backgrounds are currently excluded from the DarkSHINE simulationDarkSHINE .

### v2

@@ Abstract
This paper presents the design and optimization of a LYSO crystal electromagnetic calorimeter (ECAL) for the DarkSHINE experiment, which aims to search for dark photons as potential mediators of dark forces. The ECAL design was evaluated through comprehensive simulations, focusing on optimizing dimensions, material selection, energy distribution, and energy resolution. The ECAL configuration consists of 21$\times$21$\times$11 LYSO crystals, each measuring 2.5$\times$2.5$\times$4 cm3, arranged in a staggered layout to improve signal detection efficiency. A 4 GeV energy dynamic range was established to ensure accurate energy measurements without saturation, which is essential for background rejection and signal identification. A detailed digitization model was developed to simulate the scintillation, SiPM, and ADC behaviors, providing a more realistic representation of detector’s performance. Additionally, the study assessed radiation damage in the ECAL region, highlighting the necessity of radiation-resistant scintillators and silicon sensors.
@@ I Introduction
Dark matter (DM) Hooper:2007kb ; Clowe:2006 remains one of the most compelling mysteries in cosmology and particle physics, evident from its gravitational effects on visible matter and the cosmic microwave background. Theoretical models suggest that DM was produced through thermal processes in the early universe, with the "freeze-out" mechanismDu:2021jcj explaining its current observed density, positing a probable mass range from a few MeV to several TeV. Despite extensive searches, the specific properties and particle nature of DM remain elusive. Experiments such as XENONnTXENON:2023cxc , PandaXPandaX-4T:2021bab , CDEXWang_2021 , LUX-ZEPLINPhysRevLett.131.041002 , AMSGiovacchini:2020vxz , DAMPEKyratzis:2022gvg and LHCGiagu:2019fmp have narrowed the parameter space for weakly interacting massive particles (WIMPs)Griest:1989wd ; Ho:2012ug ; Steigman:2013yua ; Boehm:2013jpa ; Nollett:2013pwa ; Nollett:2014lwa ; Serpico:2004nm in the GeV to TeV range, yet these particles remain undetectedBillard:2021uyg .
@@
Given that high-mass DM candidate particles are yet to be detected, there is increasing interest in the sub-GeV mass range, which presents significant detection challenges due to the minimal interaction cross-sections of such light particles with ordinary matter. This has necessitated the development of innovative detection techniques distinct from traditional DM detection methods. Accelerator-based experiments are particularly promising, leveraging high-energy particle collisions to produce and detect dark photons, a hypothetical mediator between visible and dark matterHoldom:1985ag ; Foot:1991kb ; Fuyuto:2019vfe ; Choi:2020pyy ; Cheng_2022 . Facilities such as CERN’s NA64Banerjee:2019pds , LHC2024138378 , BELLE-IIPhysRevLett.130.071804 , BES-IIIPhysRevD.100.115016 ; Prasad:2019ris , and the proposed LDMXLDMX:2018cma ; LDMX:2019gvz ; Berlin:2018bsc experiment aim to explore these possibilities.
@@
Figure 1:  (a) Production of dark photons via bremsstrahlung. (b) Decay of dark photons into "invisible" modes, where the dark photon decays into dark matter particles. Fabbrichesi_2021
@@
The DarkSHINE experimentChen:2022liu ; SLi , designed to operate under the minimal dark photon model, proposes a novel approach to detecting these elusive particles. Figure 1 illustrates the conceptual mechanism of dark photon production through dark bremsstrahlung and their subsequent decay into invisible dark matter, forming the experimental basis for DarkSHINE. This experiment leverages the high-repetition-rate single-electron beam provided by the Shanghai High Repetition-Rate XFEL and Extreme Light Facility (SHINE)Wan:2022het ; Zhao:2017ood , optimized for detecting subtle signals that hint the presence of dark photons.
@@
This paper delineates the design principles of the Electromagnetic Calorimeter (ECAL) for DarkSHINE experiment, focusing on the selection of materials and structural layout. Through simulation studies, we optimized the ECAL volume, achieving a balance between signal efficiency and cost-effectiveness. We also investigated the dynamic range of energy measurement in the ECAL channels. Insights from these analyses have guided the development of potential triggering strategies for future detector. To ensure the simulated detector response closely mirrors that of a real detector, we implemented a digitization model and studied the ECAL’s energy resolution. Furthermore, we evaluated the radiation damage in ECAL region, highlighting the necessity for crystals and silicon sensors that maintain high performance in high-radiation environments.
@@ II DarkSHINE experiment and ECAL design
DarkSHINE is a fixed-target experiment that focuses on the bremsstrahlung production of dark photons and measuring their invisible decay. It utilizes a high-repetition-rate single-electron beam provided by SHINE, which is currently under construction. The beam is expected to achieve an energy of 8GeV and a repetition rate of 10MHz. This corresponds to $3\times 10^{14}$ electrons-on-target events during one year of the DarkSHINE experiment’s commissioning.
Figure 2: Schematic of the DarkSHINE detector system, illustrating the primary components used for dark photon detection. The beam enters the center of the detector along the direction from left to right in the diagram. The setup includes the magnet (red) and its supportor (blue), a tagging tracker with seven layers (equal-sized tiles with gray in the middle and green at both ends) inside the magnet, a recoil tracker with six layers (unequal-sized tiles) behind the tagging tracker, a tungsten target sandwiched between the tagging tracker and the recoil tracker, an electromagnetic calorimeter (small gray cube) behind the recoil tracker, and a hadronic calorimeter (big yellow cube) behind the electromagnetic calorimeter. Chen:2022liu
@@
The primary challenge for DarkSHINE is to use information from various detectors to minimize background contributions while preserving the dark photon signal. Figure 2 shows the detector system of DarkSHINE. It primarily consists of a tungsten target and three sub-detectors: a tracker, an electromagnetic calorimeter (ECAL), and a hadronic calorimeter (HCAL). Additionally, a single electron beam with a spatial distribution radius of 3 cm and a maximum non-uniform magnetic field of 1.5 Tesla in the tracker region were used in the simulations. The tracker measures the momentum and trajectory of electrons, consisting of a total of thirteen layers of silicon strips. Seven layers are placed in front of the target to tag the incident electron (tagging tracker), and six layers are placed behind the target (recoil tracker). The tungsten target (0.1 X0) is sandwiched between the tagging tracker and the recoil tracker. A crystal ECAL positioned behind the recoil tracker measures the deposited energy of recoil electrons and photons. The recoil electrons are defined as the state of incident electrons after interacting with or passing through the target. Then, a scintillator-steel based sampling HCAL is placed behind the ECAL to capture and veto the backgrounds, in particular the neutral hadrons and muons.
@@
To measure the invisible decay of dark photons, it is essential to precisely measure the energy of recoil electrons after each collision to determine if there is a significant energy loss. The core detector of DarkSHINE, the ECAL designed for DarkSHINE, is a homogeneous LYSO crystal calorimeter. LYSO1239590 ; Butler:2019rpu ; Kalinnikov:2023coj is an excellent choice for our crystal material, not only because of its high light yield, which is critical for electromagnetic resolution, but also due to its rapid scintillation decay time (40 ns), which is vital for handling the extremely high event rate we face. Furthermore, the ECAL, especially its central region, is exposed to a considerable radiation dose, necessitating materials with good radiation resistance. LYSO has shown exceptional durability in such conditions, establishing it as an optimal selection for DarkSHINE ECAL applications.
@@
Figure 3:  Configuration of the Electromagnetic Calorimeter (ECAL) in DarkSHINE. The ECAL consists of an array of LYSO crystals, each coupled with one SiPM for readout at the end. The schematic shows the segmentation and staggered layout of the crystals within the ECAL.
The structure of ECAL is illustrated in Figure 3. The transverse dimension of the ECAL is about 52.5$\times$52.5 cm2, with a dimension along the beam direction around 44 cm, approximately equivalent to 39 radiation lengths. The substantial depth ensures excellent energy containment in the ECAL, guaranteeing that it can absorb nearly all electromagnetic showers, preventing them from leaking into the HCAL. This is crucial because such leakage would result in vetoing dark photon signals by the HCAL. The volume of ECAL is determined by balancing the signal efficiency and crystal cost (Section IV). The ECAL is fully segmented to gather comprehensive information, enhancing its capability to discriminate dark photons. It is composed of 21$\times$21$\times$11 LYSO crystals, with each crystal measuring 2.5$\times$2.5$\times$4 cm3 covered by ESR and read out by SiPMsKLANNER201936 ; SIMON201985 . In each layer, the crystals are positioned in a uniform 21$\times$21 square pattern. To enhance detection efficacy and prevent particles from traversing the gaps, the placement of crystals in successive layers is shifted by half the transverse dimension of a crystal. This staggered structure aids in detecting recoil electron with energy loss coming along with dark photon signal. Yu:2024quo ; Xu:2024btf
III Signal and background
III.1 Software setup
@@
The Monte Carlo simulation of the ECAL is performed using the DarkSHINE Software (DSS) framework based on Geant4GEANT4:2002zbu ; Zhu:2023 . This simulation includes the entire detector setup, encompassing the crystal cubes, wrappers, silicon sensors, and support structures, alongside other components like the Tracker and HCAL. This comprehensive setup provides the necessary reference for ECAL reconstruction and the initial analysis cuts to define the appropriate phase space for ECAL analysis. Within DSS, we can simulate all Standard Model background processes involving electron-target interactions and various forms of dark photon processes, as well as their responses in the detector.
@@ III.2 Signal and background
Figure 4:  Background processes in DarkSHINE experiment, along with their relative rates, under an 8 GeV electron beam. Chen:2022liu
Figure 4 illustrates the background processes considered in DarkSHINE experiment, along with their relative rates, under an 8 GeV electron beam. In the majority of cases, the incident electron passes through the target with minimal interacting. Approximately 6.7% of electrons undergo hard bremsstrahlung, where the bremsstrahlung photons carry away more than 4 GeV from the incident electrons. These bremsstrahlung photons can then interact with the target or ECAL materials via photon-nuclear interactions, occurring with relative rates of 2.31 $\times$ 10-4 and 1.37 $\times$ 10-6, respectively. In rare instances, bremsstrahlung photons may convert into a muon pair, with relative rates of 1.63 $\times$ 10-6 in the ECAL and 1.50 $\times$ 10-8 in the target. These photons can also produce electron-positron pairs at significantly higher rates. Additionally, electron-nuclear interactions within the ECAL and target materials contribute to the background at relative rates of 3.25 $\times$ 10-6 and 5.10 $\times$ 10-7, respectively. Chen:2022liu
Figure 5:  Deposited energy of 8 GeV electrons in ECAL for dark photons and inclusive background processes. In dark photon processes, the energy deposited in ECAL is mostly less than 4 GeV, and it decreases as the mass of the dark photon increases. In contrast, inclusive background process results in significantly higher energy deposition within the ECAL.
@@
When a dark photon is produced during the electron-on-target process, most of the incident energy transfers to it, and the recoiled electron deposits its remaining energy in the ECAL. The dark photon decays into dark matter, leaving no signal in the tracker or calorimeter, causing significant energy loss in the ECAL. In Figure 5, more than 75% of the events have energy deposits in the ECAL that are less than 4 GeV for 1 MeV dark photon signals. As the dark photon mass increases, it carries more energy from the electrons, leading to larger energy loss measured in ECAL.
The inclusive process encompasses all possible Standard Model physics processes that can occur in the detector, weighted by their respective branching ratios. Compared to the dark photon signal, most events in the inclusive background tend to deposit the majority of their energy in the ECAL. Therefore, the total energy measured by the ECAL becomes a robust criterion for background rejection.
In inclusive process, when the incident electron undergoes hard bremsstrahlung or the produced bremsstrahlung photon converts into an electron-positron pair, these events generate electromagnetic showers in the ECAL, leading to significant energy deposition. However, hard processes involving production of high energy muon pairs or hadrons tend to deposit only a small fraction of energy in the ECAL, but they can be effectively rejected using the tracker and HCAL.
@@
Neutrino-producing background processes, such as Moller scattering ($e^{-}e^{-}\rightarrow e^{-}e^{-}$) followed by a charged-current quasi-elastic (CCQE) reaction ($e^{-}p\rightarrow\nu_{e}n$), neutrino pair production ($e^{-}N\rightarrow e^{-}N\nu\nu$), bremsstrahlung with CCQE, and charge-current exchange with exclusive ($e^{-}p\rightarrow\nu n\pi^{0}$), as mentioned in Izaguirre:2014bca , have been found negligible in this experiment due to their low event ratios. Therefore, these backgrounds are currently excluded from the DarkSHINE simulationChen:2022liu .